OpenStudy (anonymous):

If x+y=5,y+z=8, and z+x=9, what is the average(arithmetic mean) of x,y, and z?

OpenStudy (ash2326):

zArithmetic mean ( or Average) is \[\frac{x+y+z}{3}\] we have \[x+y=5\] \[y+z=8\] \[z+x=9\] Add all the three equations, what would you get?

OpenStudy (anonymous):

2x+2y+2z=22

OpenStudy (ash2326):

good, can you find the average now?

OpenStudy (anonymous):

ehh

OpenStudy (anonymous):

2x+2y+2z=22/6?

OpenStudy (anonymous):

ehh

OpenStudy (anonymous):

what should I do now, Son Goku?

OpenStudy (ash2326):

Average= \[\frac{x+y+z}{3} \ \ \ \ \ \ \ \ =\frac{22}{6}\] This is the mean

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

11/3?

OpenStudy (anonymous):

is that the answer, 11/3?

OpenStudy (ash2326):

yeah

OpenStudy (anonymous):

yeah!!

OpenStudy (anonymous):

you a big fan of DBZ?

OpenStudy (ash2326):

I used to like it, but haven't seen it recently :)

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

k

OpenStudy (anonymous):

ty